OpenStudy (anonymous):

s^2-6s-16=0 Solve. Would I use the quad formula?

OpenStudy (anonymous):

no it's just (s - 8) (s + 2)

OpenStudy (jlastino):

No. (s-8)(s+2) s=8 s=-2

OpenStudy (saifoo.khan):

8,-2

OpenStudy (anonymous):

(S-8)(S+2) Thus, s-8=0 s=8 s+2=0 s= -2